OpenStudy (anonymous):

if sin x cot x=1/4 wha is the value of cos x

OpenStudy (witty_roger):

cos x =1/4

OpenStudy (crabbyoldgamer):

cot x = cosx/sinx sinxcotx = sinx * cosx/sinx = cosx: sinx cancels. So cosx = 1/4

OpenStudy (anonymous):

\[\cot~x=\frac{ 1 }{ \frac{ \sin~x }{ \cos~x } }\]\[\cot~x=\frac{ \cos~x }{ \sin~x }\]substitute that into the equation\[\sin~x\color{red}{(\cot~x)}=\frac{ 1 }{ 4 }\]\[\sin~x(\frac{ \cos~x }{ \sin~x })=\frac{ 1 }{ 4 }\]\[\frac{ \cos~x ~\sin~x}{ \sin~x }=\frac{ 1 }{ 4 }\]\[\frac{ \cos~x~\cancel{\sin~x} }{ \cancel{\sin~x} }=\frac{ 1 }{ 4 }\]\[Therefore,\cos~x=?\]solve it @Justmeneedinghelp